### v1

# revtex4-1Repair the float

Abstract
We study the translocation of a flexible polymer through extended patterned pores using molecular dynamics (MD) simulations. We consider cylindrical and conical pore geometries that can be controlled by the angle of the pore apex $\alpha$. We obtained the average translocation time $\langle\tau\rangle$ for various chain lengths $N$ and the length of the pores $L_{p}$ for various values $\alpha$ and found that $\langle\tau\rangle$ scales as $\langle\tau\rangle\sim N^{\gamma}\mathcal{F}\left(L_{p}N^{\phi}\right)$ with exponents $\gamma=3.00\pm 0.05$ and $\phi=1.50\pm 0.05$ for both patterned and unpatterned pores.
I Introduction
Translocation of polymer or proteins is an ubiquitous process inside biological cells and is among one of the important processes which propel the cell functionality Alberts et al. (2002); van der Laan et al. (2007). In the last two and half decades, polymer translocation has gained considerable attention due to its technological applications such as rapid DNA sequencing, gene therapy and controlled drug delivery Salman et al. (2001); Ying et al. (2022). It is now possible to fabricate synthetic nanopores in the laboratory which can remain stable at high temperatures, external drive or change in the solvent conditions. A single strand of a biopolymer such as DNA or RNA can be electrophoretically driven through such nanopores in a controlled manner. The ion current flowing through the nanopore gets blocked when the polymer is inside the pore which gives the signature of the polymer sequence Kasianowicz et al. (1996).
The translocation of polymer through nanopores depend on various parameters, such as, the length of the polymer chain, its structure, shape and the size of the pore and pore-polymer interactions Muthukumar (2007, 2011). Due to these factors, the translocation process is a complicated system dependent process. Various coarse-grained models have been used to understand the translocation phenomenon Sung and Park (1996); Muthukumar (2011); Palyulin et al. (2014); Rowghanian and Grosberg (2011); Sakaue (2007); Kantor and Kardar (2004); Dubbeldam et al. (2007); Ikonen et al. (2012a). For example, the translocation process has been studied as a one-dimensional barrier crossing problem of the translocation coordinate through a slit (i.e., a pore of unit length). In initial studies, it was assumed that the translocation time is long enough to ensure the equilibration of the polymer conformations at every stage of the translocation process Sung and Park (1996); Muthukumar (1999, 2011). Later, a majority of experimental studies with synthetic nanopores showed that the time taken by the polymer to pass the nanopore is much less than its relaxation time, indicating that the translocation is a non-equilibrium process. The tension propagation (TP) theory proposed originally by Sakaue Sakaue (2007) for an infinite chain and subsequently modified by Ikonen et al. Ikonen et al. (2012a, b); Sarabadani et al. (2014); Sarabadani and Ala-Nissila (2018) and Dubbeldam et al. Dubbeldam et al. (2012) for finite chains has proven to be successful in explaining the non-equilibrium aspect of driven polymer translocation Chen et al. (2021). The most important quantity of interest in polymer translocation studies is the dependence of the average translocation time, $\langle\tau\rangle$, on the chain length $N$. Numerical simulations and experiments have indicated that $\langle\tau\rangle\sim N^{\beta}$. Different values of $\beta$ have been observed in different studies, suggesting that $\langle\tau\rangle$ depends on various physical parameters Milchev (2011). For a self-avoiding flexible chain, $\langle\tau\rangle$ is known to scale as $\langle\tau\rangle\sim N^{1+\nu}$, here $\nu=3/4$ in two dimensions Chuang et al. (2001); Luo et al. (2007); Sarabadani and Ala-Nissila (2018), and $\nu\approx 0.59$ in three dimensions Doi et al. (1988); Chuang et al. (2001); Kantor and Kardar (2004).
Experimental studies on polymer translocation involve pores of finite lengths in which pore-polymer interactions play a dominant role Storm et al. (2003); Wanunu and Meller (2007); Chen et al. (2004); Lan et al. (2011); Liu et al. (2012). There have been theoretical and computational studies on finite-length pores where the shape of the extended pore is cylindrical with different patterned pore-polymer interactions Luo et al. (2007, 2008); Cohen et al. (2012a, b); Katkar and Muthukumar (2018); Kumar et al. (2018), or conical Nikoofard et al. (2013, 2014); Nikoofard and Fazli (2015); Sharma et al. (2022); Sharma (2024); Upadhyay et al. (2024a, b). For cylindrical pores, it was estimated that for a fixed external drive $\langle\tau\rangle$ depends on the patterning of the pores. It was found that $\langle\tau\rangle$ is the maximum for a fully attractive pore and is minimum for a pore having repulsive exit Cohen et al. (2012b); Kumar et al. (2018). It was argued that the pattern stickiness of extended pores can be exploited to predict heteropolymer sequences with different bending rigidities to a high degree of accuracy by translocating them through multiple pores Kumar et al. (2018). For conical pores, we have one more parameter–the cone apex angle, which we represent by $\alpha$. For $\alpha=0$, we have a cylindrical pore. For a conical pore with attractive pore-polymer interactions, $\langle\tau\rangle$ was found to be a nonmonotonic function of $\alpha$ for lower driving forces, which becomes monotonic for higher forces Sharma et al. (2022); Sharma (2024). The length of the extended pore, $L_{p}$, introduces an additional length scale to the problem. Although there have been many scaling studies of $\langle\tau\rangle$ on the length of the polymer $N$, not much is discussed in the literature on the dependence of $\langle\tau\rangle$ on $L_{p}$. We try to address this in this paper.
In this paper, we extend our work on conical pores Sharma et al. (2022); Sharma (2024) to study the effect of different pore patterns on the driven-translocation process. The patterning is created along the pore length by having various sections with attractive and repulsive interactions. The entrance of the pores is always attractive, so that the polymer can enter the pore with ease. The exit from the pore can either be attractive or repulsive. The presence of an external drive along the channel often facilitates the translocation process. We present our study on translocation behavior by varying the pore length $L_{p}$ and the polymer length $N$. The length of the pore is increased in such a manner that the pore-patterning ratios remain conserved. We record the translocation time through such pores for different external drives.
The paper is organized as follows. In Sec. II, we define our model and simulation details. In Sec. III, we study the effect of the pore length, $L_{p}$, on $\langle\tau\rangle$. We also discuss the effect of pattered pores on $\langle\tau\rangle$ and establish scaling relations between $\langle\tau\rangle$, $L_{p}$ and the polymer length $N$ for extended pores. In Sec. IV, we conclude by providing a discussion of the obtained scaling exponents.
Figure 1: Top row shows a schematic diagram of pore-polymer set up for translocation in two-dimensions. The pore is 5 beads long with half apex angle $\alpha$. The pore diameter on the cis end is $d=2.25\sigma$, whereas it is $y=d+2L_{p}\tan\alpha$ on the trans end. The flexible polymer is made up of beads of diameter $\sigma=1$ connected with springs. The subsequent rows show patterned pores with different pore lengths, $L_{p}=$ 5, 10, and 15, respectively. Blue beads are attractive while yellow and green beads are repulsive in interaction with polymer chain.
II Model
We study the pore-polymer system in two-dimensions. The polymer is modeled as a bead-spring chain Grest and Kremer (1986), where the consecutive beads of diameter $\sigma$ are connected with a spring of spring constant $K$ interacting via a harmonic potential given by:
$U_{\textrm{bond}}=\frac{1}{2}K(r-r_{0})^{2},$ (1)
where, $r_{0}$ is the equilibrium distance. The non-bonded polymer beads interact via Weeks-Chandler-Andersen (WCA) potential $U_{\textrm{nb}}(r)$ of the form:
$U_{\textrm{nb}}(r)=\begin{cases}4\epsilon\Bigg{[}\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\Bigg{]}+\epsilon,&r<r_{\rm min}\\ 0,&r\geq r_{\rm min},\end{cases}$ (2)
where, $\epsilon$ is the strength of the potential. The cutoff distance $r_{\rm min}=2^{1/6}\sigma$ is set at the minimum of the potential. The pores are static and are constructed with coarse-grained beads of size $\sigma$. We study three different types of extended pores as considered in Ref. Kumar et al. (2018):
  1. (1)
    Pore A is an attractive pore. Polymer beads interact with pore beads via the attractive Lennard Jones (LJ) potential, $U_{\textrm{pp}}(r)$:
    $U_{\textrm{pp}}(r)=\begin{cases}4\epsilon_{p}\Bigg{[}\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\Bigg{]},&r<r_{c}\\ 0,&\text{otherwise},\end{cases}$ (3)
    where, $\epsilon_{p}$ is the attractive strength of the pore and $r_{c}=2.5\sigma$ is the cut-off distance.
  2. (2)
    Pore B has attractive entrances and exits separated by a repulsive section. The polymer beads interact with attractive beads of the pore via the LJ potential and with beads of the repulsive section via the WCA potential.
  3. (3)
    Pore C has an attractive entrance and a repulsive exit. The polymer beads interact with the beads of the attractive section of the pore via the LJ potential and the beads of the repulsive section via WCA potential.
In addition to the above potentials, an external pulling force, $f_{ext}$, which mimics an electrophoretic force, is applied inside the pore in the positive $x$-direction to ensure that the polymer translocates through the pore from the cis to the trans side. The force $f_{ext}$ varies along the pore extension as
$f_{ext}=\frac{f_{0}d}{(d+2x\tan\alpha)},$ (4)
here $f_{0}$ is the magnitude of the applied force at the pore opening $d$, which is equal to the constant force inside a linear channel with $\alpha=0^{\circ}$. Therefore, for a conical pore, the force decreases along its length. The schematic diagram of our coarse-grained model is shown in Fig. 1 (top row). The subsequent rows show the schematic diagram of three different pore patterns that we will study in this paper. The second row shows three different patterned pores for length $L_{p}=5$, here the pore in Fig. 1(a1) is a completely attractive pore (Pore A), while the pore in Fig. 1(b1) has an attractive entrance and an attractive exit but embedded with a repulsive bead in the middle (Pore B), and the pore in Fig. 1(c1) has an attractive entrance but a repulsive exit (Pore C). In the third and fourth rows, we show pores for lengths $L_{p}=10$, and $L_{p}=15$, respectively. The length of the pore is increased in such a manner that the ratio of pore patterning remains conserved.
The dynamics of the polymer is governed by the stochastic Langevin equation. The equation of motion for the position of $i$th monomer is given by
$m\ddot{\vec{r_{i}}}=-\eta\dot{\vec{r_{i}}}-\nabla\sum{(U_{bond}+U_{nb}+U_{pp})}+f_{ext}+\zeta,$ (5)
where $m$ and $r_{i}$ are the mass and the position of the $i^{th}$ bead of the polymer, $\eta$ is the friction coefficient of the fluid, $\zeta$ is an uncorrelated random noise that obeys the fluctuation-dissipation relation $\langle\zeta_{i}(t)\zeta_{j}(t^{\prime})\rangle=2K_{B}T\eta\delta_{ij}\delta(t-t^{\prime})$. In our model, $\epsilon$, $\sigma$ and $m$ set the units of energy, length, and mass, respectively. This sets the unit of time as $(m\sigma^{2}/\epsilon)^{1/2}$. The force value is set in the unit of $\epsilon/\sigma$ and the spring constant in the units of $\epsilon/\sigma^{2}$. Using this unit, the dimensionless parameters $k_{B}T=1$, $\eta=1$, $d=2.25$, and $K=500$ are chosen in the simulations. For every simulation data point, the results are averaged over 1500-2000 successful translocation events. MD trajectories are obtained using LAMMPSThompson et al. (2022) software.
III Result and Discussion
III.1 Effect of pore length
[…]
Figure 3: $\langle\tau\rangle$ as a function of NN (in log-log scale) for three different driving forces when the polymer is translocating from an extended cylindrical pore (i.e., $\alpha=0$) of length (a) $L_{p}=5\sigma$, and (b) $L_{p}=10\sigma$. The slopes mentioned in the plot are the values of the translocation exponent, $\beta$, defined by $\langle\tau\rangle\sim N^{\beta}$, for different driving forces.
Figure 4: $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for an extended pore ($\alpha=0$) for three different driving forces (a) 0.2, (b) 0.6, and (c) 1.0. The lines are the scaling function $\mathcal{F}(x)\sim 1/x^{p}$ with exponent $p=0.85$.
The length of the pores, $L_{p}$, plays an important role in the translocation of a polymer through a membrane channel. To study how the translocation time depends on the length of the pores, we consider a polymer of length $N=128$ and translocate it through pores of varying lengths $L_{p}\in[5\sigma$, $10\sigma]$. For each pore length, we also vary the angle of the cone apex, $\alpha$, in the range $\alpha\in[0,5]$ degrees. The magnitude of the external force is kept constant at $f_{0}=1$. The pore-polymer interaction is attractive in nature and is modeled using Eq.(3). In Fig. 2, we have plotted the average translocation time $\langle\tau\rangle$ as a function of $\alpha$ for various $L_{p}$. The plot clearly shows that, irrespective of the value of $\alpha$, $\langle\tau\rangle$ decreases with increasing pore length $L_{p}$.
For $\alpha=0^{\circ}$ the value of $\langle\tau\rangle$ obtained for $L_{p}=5\sigma$ is almost three times that for $L_{p}=10\sigma$. This can be understood as follows: At $\alpha=0^{\circ}$ the pore is cylindrical in geometry with a diameter $d=2.25\sigma$, which allows only a single-file entry and exit from the pore. For such symmetrical pores, the extent of the external force increases as the pore length is increased. For example, for a completely filled pore of length $L_{p}=10\sigma$, the number of beads experiencing force is almost twice the number of beads for that of $L_{p}=5\sigma$. Therefore, the polymer can be driven easily through a longer pore.
For $\alpha>0^{\circ}$, i.e., for conical pores, the diameter of the channel increases as one moves from the apex to the base of the pore. The diameter at the narrow entrance remains the same as that for a cylindrical pore (i.e., $d=2.25\sigma$), while the diameter of the wider end increases by an amount $2L_{p}\sin\alpha$ [i.e., $d(x)=d+2x\sin\alpha$]. Now as the area available on the wider end of the conical channel increases, the single-file nature of the translocation process is no longer true and more configurations of the polymer are possible, this gain in the conformational entropic force Nikoofard et al. (2013) assists the polymer to escape faster from the channel. This is indeed seen for shorter pore lengths, e.g., $L_{p}=5$ and 6, where $\langle\tau\rangle$ decreases on increasing $\alpha$ and saturates for $\alpha>2^{\circ}$. In contrast, for longer pore lengths, we see an opposite behavior. As the length of the pore increases, the diameter at the wider end increases proportionately. Since the external force inside the channel decreases as $f(x)\sim 1/x$ (see Eq. (4)), it also becomes weaker at the wider pore exit for longer pores. Further, since the pore is attractive in nature there is a higher possibility for the polymer to have a configuration where the two consecutive monomers of the chain occupy side-by-side position along the $y$-direction at some location $x$ in the channel, thus hindering translocation. Combining all these factors, it was observed that the total translocation time for longer pores increases as the pore apex angle $\alpha$ is increased (e.g., $L_{p}=10\sigma$ in Fig. 2).
Figure 5: Comparison of $\tau$ vs. $\alpha$ for three different pore sizes $L_{p}=$ 5, 10 and 15 for two different forces ($f_{0}=$) 0.6 and 1.0. The length(NN) of the flexible polymer chain is 128 beads.
Next, we study how $\langle\tau\rangle$ varies with the chain length $N$ for the pores of lengths $L_{p}=5\sigma$ and $10\sigma$. Let us first consider translocation through a cylindrical pore (i.e., $\alpha=0$). In Figs. 3(a) and 3(b), we have plotted $\langle\tau\rangle$ as a function of $N$ for a polymer translocating through a cylindrical pore of lengths $L_{p}=5$ and 10, respectively for various values of driving forces $f_{0}=0.2$, 0.6 and 1.0. The translocation time exponent $\beta$ is defined as $\tau\sim N^{\beta}$. Recall that, for a flexible polymer translocating from a slit, it was found that the exponent $\beta<1+\nu$, where $\nu$ is the size exponent Ikonen et al. (2012b, a). In 1- and 2-dimensions $\nu=1.0$ and $0.75$, respectively. It is known that the value of exponent $\beta$ depends on many factors such as chain length, chain flexibility, driving force, friction, and pore structure. This is also observed in our simulations. For a smaller driving force $f_{0}=0.2$, we found that the numerical value of the exponent for pore length $L_{p}=5$ is $\beta=1.32$, which is smaller than $1+\nu=1.75$. On increasing the pore length to $L_{p}=10$, the value of the exponent becomes $\beta=1.68$ but still less than 1.75. On increasing the driving force to $f_{0}=0.6$, the translocation exponent for $L_{p}=5$ is $\beta=1.60<1.75$. However, for for pore length $L_{p}=10$, the exponent becomes $\beta=1.87>1.75$. For a driving force $f_{0}=1.0$, the value of the translocation exponent for pore lengths $L_{p}=5$ and 10 are 1.80 and 1.95, respectively, both are greater than 1.75 but less than 2.
We have seen that the translocation exponent $\beta$ varies with the length of the extended pore. We now seek to see whether the average translocation time $\langle\tau\rangle$ for different pore lengths $L_{p}$ and the polymer chain lengths $N$ follow some scaling. To check this, we take the scaling form for $\langle\tau\rangle$:
$\langle\tau\rangle\sim N^{\gamma}\mathcal{F}\left(L_{p}N^{\phi}\right),$ (6)
where $\gamma$ and $\phi$ are exponents, and $\mathcal{F}$ is the scaling function. In Fig. 4, we have plotted $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for pore lengths $L_{p}$ varying from $5$ to $10$ and chain lengths varying from $N=160$ to 512. A good data collapse is obtained for exponent values $\gamma=3.0\pm 0.05$ and $\phi=1.50\pm 0.05$ for three driven force values ($f_{0}=0.2$, 0.6, and 1.0) discussed in this paper. The scaling function follows a power law, $\mathcal{F}(x)\sim 1/x^{p}$, with a consistent exponent $p\sim 0.85$ for each of the three forces $f_{0}=0.2$, 0.6, and 1.0.
Figure 6: $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for patterned pores: Pore B and Pore C for $\alpha=0$ and three different pore lengths $L_{p}=5$, 10 and 15 and various chain lengths $N$ as indicated in the plots. Plots (a), (b) are for force $f_{0}=0.6$ and plots (c), (d) are for force $f_{0}=1.0$, respectively. The solid line is the scaling function of the form ${\mathcal{H}}(x)\sim 1/x^{p}$ with exponent $p=0.8$.
Figure 7: $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for patterned pores (a) Pore A, (b) Pore B and (c) Pore C for $\alpha=3^{\circ}$ and $f_{0}=1.0$. Three different pore lengths used are $L_{p}=5$, 10 and 15 and various chain lengths $N$ as indicated in the plots. The solid line is the scaling function of the form ${\mathcal{H}}(x)\sim 1/x^{p}$ with exponent $p=0.6$.
III.2 Effect of pore patterning
In this section, we discuss the effect of pore patterning (shown in Fig. 1) on $\tau$ for a polymer of length $N=128$. The translocation of flexible and semifexible polymers through cylindrical and conical pores has been studied in the past Cohen et al. (2012a); Kumar et al. (2018); Upadhyay et al. (2024a, b). It was observed that the translocation time and its standard deviation can be used to sequence the polymer by translocating the polymer multiple times through different patterned pores. Our interest in this paper is to explore the dependence of $\tau$ on the pore length $L_{p}$ of various patterned pores. We increase the pore length in such a manner that the ratio of pore patterning remains conserved. As seen in the second, third and fourth rows of Fig. 1, the ratio of repulsive pore beads (yellow bead) to attractive pore beads (blue) remains the same for three pores with pore lengths $L_{p}=5$, 10, and 15.
In Fig. 5, we have plotted the total translocation time $\tau$ for a flexible polymer of length $N=128$, as a function of the angle of the pore apex $\alpha$ for three different pore lengths $L_{p}=5$, 10, and 15 at two different values of the driving force $f_{0}=0.6$ and 1.0. The figure clearly shows that the behavior of $\tau$ strongly depends on the type of pore and its length. For shorter pore lengths (e.g. $L_{p}=5$), we observe that for both force values, $\tau$ decreases with $\alpha$ and becomes saturated. This is true for the three types of pores studied in this paper (see Figs. 5(a) and 5(d)). Upon comparing $\tau$ among these pores, we observed that the pore with a repulsive exit (shown by triangles) favors faster translocation at all angles. This is followed by the pore with an attractive entrance and exit with a repulsive bead in the middle (shown by circles), whereas the translocation is the slowest for a fully attractive pore (shown by squares). For intermediate-length pores (that is, $L_{p}=10$), the translocation time $\tau$ first increases with $\alpha$ for the force values $f_{0}=0.6$ and 1.0. For a fully attractive pore (see Fig. 1(a2)) it reaches a maximum at about $\alpha\approx 3^{\circ}$ and then decreases as $\alpha$ increases further. For pore types with repulsive beads, either in the middle (see Fig. 1(b2)) the middle portion or at the exit (see Fig. 1(c2)), the value of $\tau$ becomes saturated around $\alpha\approx 3^{\circ}$ (see Figs. 5(b) and 5(e)). The above feature is more pronounced for lower force values $f_{0}=0.6$. When the pore length increases further (e.g. $L_{p}=15$), the translocation time first decreases with $\alpha$. It reaches a minimum value at some value of $\alpha$, which is different for different types of pores, and then $\tau$ increases as $\alpha$ increases further (see Figs. 5(c) and 5(f)).
To obtain the scaling of the averaged total translocation time with pore length $L_{p}$ and chain length $N$ for patterned pores, we again use the scaling relation of the form:
$\langle\tau\rangle\sim N^{\gamma}\mathcal{H}\left(L_{p}N^{\phi}\right),$ (7)
where $\phi$ and $\gamma$ are exponents, and $\mathcal{H}$ is the scaling function. We observe that when $\langle\tau\rangle/N^{\gamma}$ is plotted as a function of $L_{p}N^{\phi}$, for various pore lengths $L_{p}$ and chain lengths $N$, the data collapse to a scaling curve $\mathcal{H}(x)\sim 1/x^{q}$. The exponents $\gamma$ and $\phi$ are remarkably robust, showing no dependence on whether the pore is cylindrical (i.e. $\alpha=0$) or conical ($\alpha\geq 0$) in shape. Adjusting the proportionality constant in the scaling function $\mathcal{H}$, we consistently obtain the exponent $q\sim 0.8$ for the cylindrical pore and $q\sim 0.6$ for the conical pore, respectively. Furthermore, the exponents do not vary much with $f_{0}$. The data collapses for $\alpha=0$ and $\alpha=3$ are shown in Figs. 6 and  7, respectively.
IV Conclusion
We have studied the effect of polymer length, pore length, and pore patterning on the total translocation time $\tau$ for cylindrical pores (i.e. $\alpha=0$) and conical pores with various apex angles $\alpha$. For a polymer of fixed length $N=128$ (see Fig. 2), we found that the behavior of the total translocation time $\tau$ changes from a decreasing function of $\alpha$ to a strictly increasing function of $\alpha$ when the pore length increases from $5\sigma$ to $10\sigma$. The two opposite translocation behaviors seen for the shorter and the longer pore lengths are nonscalable, showing the nonequilibrium nature of the driven polymer translocation process. We also observe that for a fixed angle of the pore apex $\alpha$, the translocation time is a decreasing function of the pore length ($L_{p}$). Because the extent of the external driving force is larger for longer pore lengths, the driven translocation process is favored for longer pores.
We find that, the average total translocation time $\langle\tau\rangle$ scales with the polymer lengths $N$ as $\langle\tau\rangle\sim N^{\beta}$, where the translocation exponent $\beta$ depends on the pore length $L_{p}$ as well as the driving force. We found that for smaller driving force $f=0.2$ the value of the exponent is $\beta<1+\nu=1.7$ for pore lengths $L_{p}=5$ and 10. However, for large driving force $f=1.0$, the exponent $1.75<\beta<2.0$ for both the pore lengths. For single bead pores, Kantor and Kardar argued the upper bound of the exponent to be 1.75Kantor and Kardar (2004). We present a new set of exponents for extended pore, which is different from single bead exponent. A plausible explanation could be that by increasing the pore length or the driving force reduces the embedded dimension for the polymer from two- to one and hence the upper bound for the translocation exponent for an extended pore should be 2.
We also observed that the translocation time, $\langle\tau\rangle$, scales with the pore length $L_{p}$ and the polymer length $N$ according to the relation $\langle\tau\rangle\sim N^{\gamma}\mathcal{F}(L_{p}N^{\phi}$), where $\gamma=3.0\pm 0.05$ and $\phi=1.5$ for an attractive pore across all driven forces used in our study. The scaling function $\mathcal{F}(x)$ was found to follow the form $\mathcal{F}(x)\sim 1/x^{p}$, with $p=0.85$, which remains robust for all force values, including $f_{0}=0.2$, 0.6, and 1.0. Notably, the scaling exponents were consistent for both cylindrical and conical patterned pores. These findings demonstrate that the translocation process is scale-invariant, suggesting the existence of universal physical laws governing the dynamics of polymer translocation across different scales.
V Acknowledgement
The authors thank the IISER Mohali computing facility for providing the necessary resources and support for this work.